Three friends ate dinner at a restaurant. The friends decided to add a 15% tip and then split the bill evenly. Each friend paid ten dollars and thirty five cents. What was the total bill for dinner before tip?
Answer by mananth(16949)

each of them paid 10.35
total amount paid = 10.35 * 3= 31.05
less 15% of 31.05 will be the bill
0.15*31.05=4.66 was the tip
31.05 - 4.66= 26.39 was the bill before tip